OpenStudy (anonymous):

solve for y: please show all steps 3x-4y=10

OpenStudy (anonymous):

3x-4y=10 -3x -3x (subtract 3x) -4y = 10 - 3x /-4 /-4 (divide by -4) y = (10 - 3x) / -4